OpenStudy (anonymous):

What is the relationship between the force of gravity and the motion of objects in our solar system.

OpenStudy (irishboy123):

in a nutshell, \(F = \dfrac{GMm}{r^2} = \dfrac{m v^2 }{r}\), approximating the orbits as circles.

OpenStudy (anonymous):

@irishboy123 your answer has nothing to do with my question

OpenStudy (irishboy123):

lol!! it is the answer in a nutshell, i think. maybe you were looking for something more wordy?!